OpenStudy (anonymous):

Which property could be used to find the missing number? –4 • (–6.2 + 5) = –4 • + (–4) • 5 A. distributive property B. commutative property C. associative property

OpenStudy (igreen):

What do you think?

OpenStudy (anonymous):

b

OpenStudy (igreen):

Distributive property: \(\sf x(a + b) \rightarrow x \cdot a + x \cdot b\) Commutative property: \(\sf a + b \rightarrow b + a~~~~~\large OR ~~~~~ \normalsize a \cdot b \rightarrow b \cdot a\) Associative property: \(\sf (a + b) + c\rightarrow a + (b + c)~~~~~\large OR ~~~~~ \normalsize (a \cdot b) \cdot c \rightarrow a \cdot (b \cdot c)\)

OpenStudy (igreen):

Commutative property is incorrect.

OpenStudy (anonymous):

a then

OpenStudy (anonymous):

right

OpenStudy (anonymous):

A

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@iGreen is it a